OpenStudy (anonymous):

hey..please help on this one. A function f is continuous at 1 , and it is given that lim ->2 , f(3x-5) - 2/(x-2) = 6 .now I am supposed to find the function f(x). Tried but couldnt figure out. Is it even doable?

OpenStudy (anonymous):

alright, we know that lim(xapproaches2) f(3x-5)-2/(x-2)=6, then f(6-5)-2/(2-2)=6 then f(1)=6+2/0=infinity, f(1)=infinity; there are a number of functions which satisfy this coordinate, -ln(x-1) comes to mind.

OpenStudy (anonymous):

thanks for answering... I aint sure if i got this right but as you have mentioned that f(1) -> infinity but initially they have given f is continuous at 1 ...so isnt it contradictory...please help

OpenStudy (anonymous):

continuity and value are unrelated, as long as the function is differentiatable at that point it can be concluded to be continuous there as well, since d/dx -ln(x-1)=-1/(x-1)= g(x) and g(1)=infinity the original function is continuous just with a vertical slope.

OpenStudy (anonymous):

thank you coureges...its making sense now ..although its still a bit diificult for my inner to acccept... one more question..how do i assign you a medal??

OpenStudy (anonymous):

I have absolute no idea.

OpenStudy (anonymous):

okay..let me play around the website..

OpenStudy (anonymous):

anyways thanks..